OpenStudy (xochital):

Solve the inequality 4x - 5 < 2x + 11 x > 8 x > -8 x < 8 x < -8

OpenStudy (ja1):

wait a second, hmm, im wrong my bad you have to change the < symbol to = so it would be 4x-5 = 2x + 11

OpenStudy (ja1):

now combine like terms

OpenStudy (xochital):

so it would be 6x? or 2x?

OpenStudy (ja1):

2x you have to subtract 2x from both sides

OpenStudy (ja1):

so now you have 2x - 5 = 11

OpenStudy (ja1):

\[2x - 5 = 11\]

OpenStudy (ja1):

so now you would have to isolate 2x

OpenStudy (ja1):

so you...

OpenStudy (xochital):

add 5 to both sides..

OpenStudy (ja1):

yes since 5 is a negative

OpenStudy (ja1):

\[2x = 16\]

OpenStudy (xochital):

so x = 8

OpenStudy (ja1):

not quite the math is correct but you need the < sign back but good your right

OpenStudy (ja1):

but good job hope i helped

OpenStudy (xochital):

alright! thank you so much :)

OpenStudy (ja1):

no prob :)